# Temporal Scaling Law for Large Language Models

### v2

@@ 4 Use Case #1 of Temporal Scaling Law: Hyperparameter Selection
Table 4: ID-Val perplexity of models pre-trained on different weighting strategies. “Head” represents “Head suppression”, etc. Bold represents the best result.
Specifically, recognizing that smaller models can still identify some effective, though suboptimal, hyperparameters, we employ a two-stage retrieval-rerank pipeline to select the best data mixture proportion for training a 468M/1.2B model. In the retrieval stage, following the prior art (Xie et al., 2023a), we use a small-scale model to locate a group of fairly good data mixture proportions. In the rerank stage, we train all located data proportions on the target LLM for acceptable training tokens, and locate the best one by predicting their final losses with our temporal scaling law.
Experiment Setup. In the retrieval stage, we conduct grid search on a 58M model and locate the Top-5 candidates. In the rerank stage, we train each proportion for 10B tokens before making the test loss prediction. To compare different data mixture proportions on the 468M/1.2B models, we primarily use the perplexity metric on the ID-Val and the OOD-Val. We also follow (Touvron et al., 2023a; Rae et al., 2021; Su et al., 2023) and employ 7 popular common sense reasoning benchmarks, including BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), OpenBookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), StoryCloze (Mostafazadeh et al., 2016), and Winogrande (Sakaguchi et al., 2021) for evaluation. See Appendix for more implementation details.
Experiment Result. We compare the data mixture proportion selected by our retrieval-rerank pipeline with baselines (1) the Top-1 selected directly on the much smaller model and (2) the Top-1 selected by comparing the real test losses after training 10B tokens on 468M/1.2B model. We find that both (1) and (2) make the same choice from the Top-5 candidates, and our pipeline locates a different choice from them. As shown in Tab. 2, our choice achieves lower perplexity across all settings after completing pre-training (i.e., 400B tokens). Additionally, as shown in Tab. 3, our choice yields superior benchmark performance in 12 of 14 metrics. Consequently, while optimal hyperparameter selected on small models may not transfer to large models, our temporal scaling law allows for a more precise selection of training hyperparameters based on the candidates filtered out by small models.
5 Use Case #2 of Temporal Scaling Law: Revisit the Training Strategy
The temporal scaling law has afforded the possibility to study LLM pre-training in a finer granularity. As noted in Sec. 3.3, tokens on different positions possess a fundamental bias in learning difficulty. Specifically, head tokens are generally harder to predict than tail tokens. However, as suggested by Eqs. 6 and 8, $a_{0}$ and $a_{1}$ remains constant after the separation point. Therefore, for $N>N_{sep}$:
which is unrelated to token position $i$, suggesting that LLMs learn equally on different token positions. Considering the uniform learning dynamics across token positions, we hypothesize that the default strategy for training generative language models, in which losses on tokens in all positions are simply averaged, is an effective solution.
To validate the hypothesis, we conduct LLM pre-training with three position-based weighting strategies: Head suppression, Body suppression, and Tail suppression. These strategies are attached only after the $N_{sep}$ point. As shown in Tab. 4, despite being weighted on different token positions, all suppression strategies acquire comparable or slightly inferior results to the default practice, in which losses are simply averaged among positions. It indicates that weighting different tokens by position during LLM pre-training probably cannot yield better results than the default practice, further demonstrating that it is unnecessary to re-weight tokens by their positions during LLM pre-training, though different positions possess fundamental bias in learning difficulty. See Appendix for more implementation and analysis details.
6 Conclusion
Our research introduces the novel concept of Temporal Scaling Law for Large Language Models (LLMs), studying how the loss of an LLM evolve as the training steps scale up. We analyze the loss patterns on different token positions and discovered that these patterns conform to a dynamic hyperbolic-law. By studying the temporal evolution of the dynamic hyperbolic-law, we could properly fit and precisely predict the evolution of the LLM’s test loss, marking a significant improvement over baseline methods. Such capability is crucial for numerous possible applications. For hyperparameter selection, we provide a retrieval-rerank pipeline, in which our temporal scaling law can be paired with hyperparameter tuning methods for better hyperparameter selection based directly on the target LLM’s performance. For understanding LLM pre-training, we provide evidence supporting that LLMs uniformly optimize the losses across different token positions after an early training period.
7 Limitations.
Our work termed Temporal Scaling Law, pioneers the modeling of LLM pre-training from a temporal perspective. Although our temporal scaling law exhibits significant potential for both LLM research and application, limitations still exist. First, our temporal scaling law is derived from the mainstream structure of LLMs, the GPT-based decoder transformers. The applicability of our findings to other architectures, such as Mixture of Experts (MoE) models, remains to be explored. Furthermore, our research primarily focuses on the pre-training stage. The temporal patterns in other scenarios, such as transfer learning, are not covered. We leave further investigations of the above topics to our future research.
Appendix
Appendix A Complete Fitting Results
As a supplement to Fig. 1, we provide fitting results of the dynamic hyperbolic-law on the OOD-Val for both the 468M and 1.2B models in Fig. 4. As a supplement to Fig. 2 in the article, we provide fitting results on both the ID-Val and the OOD-Val for both the 468M and 1.2B model scales in Fig. 5. Across different settings on model scales and validation set distributions, our temporal scaling law is able to depict the general trend of parameter evolution and reduce the influence of fluctuations as much as possible. By aggregating the temporal patterns on each token position through Eq. 10, our temporal scaling law is effective for modeling the overall training trajectory, reaching $R^{2}>0.99$ for all settings, as shown in Tab. 1 in the article.
Appendix B Further Experiment Details
In this section, we provide further experiment details towards the application cases in Sec. 4 and Sec. 5 in the article.
B.1 Additional General Information
Due to the page limit, we omitted some details for the general settings of our experiments. They are thoroughly described below for reproducibility.
Evaluation Pipeline. For all benchmark evaluations, we utilize the open-source LLM evaluation tool lm-evaluation-harness [^2] https://github.com/EleutherAI/lm-evaluation-harness (Gao et al., 2023), following (Su et al., 2023; Biderman et al., 2023). For all numerical results, we report the average result tested on the last 5 checkpoints.
License for Scientific Artifacts. The Pile (Gao et al., 2021) is subject to the MIT license [^3] https://arxiv.org/pdf/2201.07311. The C4 dataset (Raffel et al., 2020) is licensed under Open Data Commons License Attribution family [^4] https://huggingface.co/datasets/allenai/c4. The evaluation benchmarks (Gao et al., 2023) are subject to the MIT license. All usages of scientific artifacts in this paper obey the corresponding licenses.
Hyperparameters of Models Used. We report the details of model hyperparameters in Tab. 5.
B.2 More Experiment Setting Details for Hyperparameter Selection
@@
Due to the page limit, we omitted some details for the experiment settings in Sec. 4. They are thoroughly described below for reproducibility.
@@
58M Model Construction. The architecture of the 58M model is identical to the 70M parameter model outlined in Biderman et al. (2023). The differences between model parameters can be ascribed to different vocabulary sizes and different activation functions (i.e., SwiGLU v.s. GeLU).
@@
Proportion Candidate Generation. In the retrieval stage, we generate a group of proportion candidates by conducting grid search on domain weights in the Pile dataset Gao et al. (2021). Specifically, we first calculate the original domain weight according to the original epoch settings noted by Gao et al. (2021) on our applied 32k tokenizer. Note that this is also the domain weight applied in Sec. 3.2. For the grid search, we adopt a simple but popular pipeline by selecting a domain, modifying its domain weight by a random value $s\in[-0.05,0.05]$, and normalizing the domain weight in other domains.
@@
B.3 More Observations for Token Position Loss
@@
In Sec. 5, we have stressed that a fundamental bias in learning difficulty based on token position exists. Specifically, head tokens (with shorter context) are generally harder to predict than tail tokens (with longer context), due to the increased uncertainty caused by limited contextual information available for earlier tokens in a sequence. Surprisingly, our temporal scaling law suggests that LLMs learn equally on different token positions after an early training period (as shown by Eq. 17), despite the learning difficulty bias.
@@
Observation for Actual Loss Decrease on Positions. To validate this observation, we plot the actual loss decrease observed during training along different token positions in Fig. 6. As shown in Fig. 6, in the early training period of “from 20B to 40B tokens”, the head tokens suffer from less loss decrease due to a higher learning difficulty. However, after the early training period, in a later period of ”from 140B to 160B tokens” that the separation point conditions in Eq. 7 are already fulfilled, test loss decrease among all token positions tends to be uniform across all settings. Therefore, we can infer from this observation that the learning dynamics derived from our temporal scaling law are authentically presented in LLM pre-training.
@@
Details for Weighting Strategies. In Sec. 5, we hypothesize that simply averaging the losses on all tokens is an effective solution for LLM pre-training. To validate the hypothesis, we have conducted LLM pre-training with three simple position-based weighting strategies. The implementation details of these strategies are described in Tab. 7. Note that for the Suppression strategies, we normalize the weighted losses to ensure the average weight for each token position is 1.0, and thus make the corresponding average loss comparable with the default practice. We apply the pre-training settings as in Sec. 3.2 for comparing different position-based weighting strategies.
@@
Benchmark Performance Observations. In this paper, we have shown the effectiveness of the default practice from the perplexity perspective. To further validate that the default practice actually trains a competitive model compared to the weighting strategies, we test the model performance on the Common Sense Reasoning benchmark described in Sec. 4, and report average model performance on 0-shot, 1-shot, and 5-shot settings. As shown in Tab. 8, all position-based weighting strategies acquire comparable or slightly inferior average results to the default practice, in which no weighting strategies are attached. On individual tasks, the default practice even achieves top-2 accuracies among 11 of 14 settings, surpassing all weighting variants. This indicates that weighting different tokens by position during LLM pre-training probably cannot yield better results than the default training strategy, further demonstrating that it is unnecessary to re-weight tokens by their positions during LLM pre-training.

### v3

@@ E.2 More Results of Temporal Scaling Law
Table 10: ID-Val perplexity of models pretrained on different weighting strategies. “Head” represents “Head suppression”, etc. Bold represents the best result.
E.3 Fitting Results on More Datasets
In the article, we have proposed the temporal scaling law to precisely fit the temporal evolution of test loss during LLM pretraining. To better demonstrate the validity of our temporal scaling law, we also validate the fitting results of the 9.8M and 58M models on the LAMBADA Paperno et al. (2016) and the WikiText Merity et al. (2017) datasets. Specifically, we use the validation and the test splits for the LAMBADA dataset, and use the wikitext-2-v1 domain for the WikiText dataset. The results are shown in Tab. 8. We also provide the fitting results for the larger 468M, 1.2B, and 6.7B models. As shown in Tab. 8, all results achieve $R^{2}>0.98$, well demonstrating the effectiveness of our temporal scaling law.
BoolQ HellaSwag OpenBookQA PIQA SIQA StoryCloze Winogrande
468M Default Practice 52.28 45.39 31.87 68.64 44.04 64.24 55.33
Head Suppression 52.05 45.60 31.80 68.50 44.13 64.71 55.38
Body Suppression 51.98 45.12 31.07 68.66 43.76 64.41 55.64
Tail Suppression 52.91 45.28 31.67 68.05 44.03 64.64 55.62
1.2B Default Practice 61.42 54.07 34.20 72.00 45.87 68.63 58.64
Head Suppression 59.96 54.49 34.33 71.75 45.36 68.39 58.30
Body Suppression 60.57 53.75 34.80 72.05 44.75 68.20 57.72
Tail Suppression 61.57 54.02 35.53 71.49 45.63 68.61 58.27
Table 11: Average performance (0/1/5-shot) on common sense reasoning benchmarks of models pre-trained under different weighting strategies. Bold and underline represent the best and the second-best averaged results, correspondingly.
Appendix F Generalization to More Model Structures
To demonstrate the generalizability of the Temporal Scaling Law to other model structures, we pre-trained a 468M model based on the GPTNeoXForCausalLM Andonian et al. (2023) architecture and a 1.4B MoE model with 240M activated parameters based on the LLaMA architecture.
Experimental Setup. The 468M GPTNeoXForCausalLM model shares the same hyperparameters with the 468M LLaMA-based model listed in Tab. 5. The 1.4B MoE model is built based on the LLaMA architecture with 16 total experts and adopts a Top-2 activation strategy, following the common setting in various MoE research Su et al. (2024a, b); Huang et al. (2024). Both models are pre-trained on the Pile dataset used in the main article. Specifically, the 468M GPTNeoXForCausalLM model is pre-trained with 200B tokens, and the 1.4B MoE model is pre-trained with 100B tokens. Detailed settings are listed in Tab. 6.
Experiment Results. We present the fitting results on both ID-Val and OOD-Val of the two pre-trained models in Fig. 4. The proposed temporal scaling law successfully genralizes to those architectures, capturing the primary patterns of parameter evolution for the GPT-NeoX and the MoE models, well demonstrating its generalizability.
Appendix G Test Loss Prediction Results on OOD-Val
In Fig. 3, we presented the test loss prediction results on the ID-Val for the scaled-up larger models. Additionally, we present the prediction results on the OOD-Val in Fig. 5. Same with results on the ID-Val, our temporal scaling law consistently generates reliable results over the comparing baselines.
Appendix H More Details for Use Case #1: Hyperparameter Selection
@@
Due to the page limit, we omitted some details for the use case #1 in Section 4 in the article. They are thoroughly described below for reproducibility.
@@
58M Model Architecture. The architecture of the 58M model is identical to the 70M parameter model outlined in Biderman et al. (2023). The differences between model parameters can be ascribed to different vocabulary sizes and different activation functions (i.e., SwiGLU v.s. GeLU).
@@
Proportion Candidate Generation. In the retrieval stage, we generate a group of proportion candidates by conducting grid search on domain weights in the Pile dataset Gao et al. (2021). Specifically, we first calculate the original domain weight according to the original epoch settings noted by Gao et al. (2021) on our applied 32k tokenizer. Note that this is also the domain weight applied in Section 3.2 of the article. For the grid search, we adopt a simple but popular pipeline by selecting a domain, modifying its domain weight by a random value $s\in[-0.05,0.05]$, and normalizing the domain weight in other domains.
@@
Appendix I More Details for Use Case #2: Revisit the Training Strategy
@@
In Section 5 of the article, we have stressed that a fundamental bias in learning difficulty based on token position exists. Specifically, head tokens (with shorter context) are generally harder to predict than tail tokens (with longer context), due to the increased uncertainty caused by limited contextual information available for earlier tokens in a sequence. Surprisingly, our temporal scaling law suggests that LLMs learn equally on different token positions after an early training period (as shown by Equation 17 in the article), despite the learning difficulty bias.
@@
Observation for Actual Loss Decrease on Positions. We use the larger 468M and the 1.2B models to validate our suggestion. To validate this observation, we plot the actual loss decrease pattern observed during training along different token positions in Fig. 11. As shown in Fig. 11, in the early training period of “from 20B to 40B tokens”, the head tokens suffer from less loss decrease due to a higher learning difficulty. However, after the early training period, in a later period of ”from 140B to 160B tokens” that the separation point conditions in Equation 7 in the article are already fulfilled, test loss decrease among all token positions tends to be uniform across all settings. Therefore, we can infer from this observation that the learning dynamics derived from our temporal scaling law are authentically presented in LLM pre-training.
@@
Insights for Weighting Strategies. We hypothesize that the default strategy for training generative language models, in which losses on tokens in all positions are simply averaged, is an effective solution. To validate the hypothesis, we conduct LLM pretraining on the larger 468M and the 1.2B models with three position-based weighting strategies: Head suppression, Body suppression, and Tail suppression, applied only after the $N_{sep}$ point. The implementation details of these strategies are described in Tab. 9. Note that for the Suppression strategies, we normalize the weighted losses to ensure the average weight for each token position is 1.0, and thus make the corresponding average loss comparable with the default practice. We apply the pre-training settings as in Section 3.2 of the article for comparing different position-based weighting strategies. As shown in Tab. 10, despite being weighted on different token positions, weighting different tokens by position during LLM pretraining probably cannot yield better results than the default practice, though different positions possess fundamental bias in learning difficulty.
@@
To further validate that the default practice actually trains a competitive model compared to the weighting strategies, we test the model performance on the Common Sense Reasoning benchmark described in Section 4 of the article, and report average model performance on 0-shot, 1-shot, and 5-shot settings. As shown in Tab. 11, all position-based weighting strategies acquire comparable or slightly inferior average results to the default practice, in which no weighting strategies are attached. On individual tasks, the default practice even achieves top-2 accuracies among 11 of 14 settings, surpassing all weighting variants. This indicates that weighting different tokens by position during LLM pre-training probably cannot yield better results than the default training strategy, further demonstrating that it is unnecessary to re-weight tokens by their positions during LLM pre-training.
Figure 6: More dynamic hyperbolic-law fitting results for the 9.8M model after training for 100B, 200B, 300B, and 400B tokens on both the ID-Val and OOD-Val.
Figure 7: More dynamic hyperbolic-law fitting results for the 58M model after training for 100B, 200B, 300B, and 400B tokens on both the ID-Val and OOD-Val.
Figure 8: Temporal scaling law fitting results for the 58M model on the OOD-Val.
Figure 9: More temporal scaling law fitting results for the larger 468M and the 1.2B models on both the ID-Val and OOD-Val.
Figure 10: More temporal scaling law fitting results for the 6.7B model on both the ID-Val and OOD-Val.
Figure 11: Test loss decrease on different token positions in the given training period for the 468M and the 1.2B models. “$\Delta_{20B-40B}$” means the decrease of loss at each token position from “being trained with 20B tokens” to “being trained with 40B tokens” etc. After an early training period, the loss decrease tends to be uniform across all token positions (i.e., $\Delta_{140B-160B}$).